A satellite moves in an elliptical orbit with the Earth at one focus. At the perigee its
speed is v and its distance from the centre of the Earth is R . If the eccentricity of its
orbit is 0.5, calculate its speed at the apogee.
1
Expert's answer
2019-03-28T05:31:10-0400
"R_p=R"

"e=0.5=\frac{R_a-R}{R_a+R}"

"e=0.5=\frac{Rv^2}{GM}-1"

"GM=\frac{2}{3}Rv^2"

"R_a=3R"

"v_a=\sqrt{\frac{2GMR}{R_a(R_a+R)}}"

"v_a=\sqrt{\frac{2\frac{2}{3}Rv^2R}{3R(3R+R)}}=\frac{v}{3}"